Let's consider a circle in the plane with radius r and on it a rope AB at a distance r/2 from the center. Let C be a point of the minor arc AB. So the angle ACB...


A. measures 60^0


B. measures 120^0


C. has a measure that depends on the position of C


1
Expert's answer
2021-08-12T12:48:16-0400

Let "SO" be the distance from the line segment "AB" to the center of the circle "O", then "SO \perp AB", hence, by the Pythagorean theorem, "AB=2AS=\sqrt{AO^2-SO^2}=\sqrt{r^2-\frac{1}{4}r^2}=r\sqrt{3}".

By the cosine theorem

"⌒ACB=\angle{AOB}=\arccos{\bigg(\dfrac{AO^2+BO^2-AB^2}{2×AO×BO}\bigg)}=\arccos{\bigg(\dfrac{-r^2}{2r^2}\bigg)}=\arccos{(-0.5)}=120\degree"

Answer is B.